# Integrated Open-Source Framework for Simulation of Transcatheter Pulmonary Valves in Native Right Ventricular Outflow Tracts

### v2

@@ 1 ABBREVIATIONS
2D
Two Dimensional
3D
Three Dimensional
CHD
Congenital Heart Disease
CTA
Computed Tomography Angiography
DICOM
Digital Imaging and Communications in Medicine
FE
Finite Element
HGO
Holzapfel Gasser Ogden
RVOT
Right Ventricular Outflow Tract
RV-PA
Right Ventricle to Pulmonary Artery
STL
Stereolithography
ToF
Tetralogy of Fallot
TPV
Transcatheter Pulmonary Valve
TPVR
Transcatheter Pulmonary Valve Replacement
@@ 2 INTRODUCTION
Tetralogy of Fallot (ToF) is a form of congenital heart disease (CHD) characterized by pulmonary valve stenosis, right ventricular hypertrophy, a ventricular septal defect, and an overriding aorta. ToF is the most common form of cyanotic CHD, with a prevalence of about 5 out of every 10,000 live births [1]. In most cases of ToF, a surgical patch is placed to enlarge the pulmonary valve annulus (transannular patch). This patch relieves the obstruction to blood flow in infancy but results in loss of competence of the pulmonary valve. While initially tolerated, over time, severe regurgitation of the pulmonary valve results in right ventricular dilation, ventricular failure, and life-threatening arrhythmias [2, 3, 4, 5]. Surgical implantation of a bioprosthetic valve has been used to remedy pulmonary insufficiency and related complications, but has the associated morbidity of open heart surgery on cardiopulmonary bypass [6, 7, 8, 9, 10].
@@
Over the last two decades, transcatheter pulmonary valve replacement (TPVR) has evolved to avoid the risks of open-heart surgery [9, 10, 11]. The first transcatheter pulmonary valve (TPV) devices available, the Melody valve (Medtronic, Minneapolis, MN) and Sapien valve (Edwards Lifesciences, Irvine, CA), were balloon-expandable bioprosthetic valves approved only for implantation in the small percentage of patients who received surgically placed cylindrical right ventricle-to-pulmonary artery (RV-PA) conduits as part of their initial repair [9, 10, 12]. However, the vast majority of patients with repaired ToF have surgically patched native outflow tracts, which demonstrate significant anatomic variability [13, 14]. As such, nearly 80% of patients with ToF are not candidates for traditional balloon expandable TPV therapies [7, 9, 10, 11, 12, 13, 15].
@@
In order to meet this need, self-expanding TPV systems that can conform to a wide range of heterogeneous native RVOTs have been developed. In contrast to balloon-expandable TPVs, self-expanding TPVs consist of a spring-like Nitinol frame enclosing a tissue valve [16, 17, 18], which can expand to conform to the vessel and facilitate device retention. However, precise fit of the device in a given individual is critical; if the device is dislodged, it can catastrophically occlude the pulmonary outflow. Further, if the proximal and distal regions of the device do not effectively seal the pulmonary artery, a paravalvular leak or device thrombosis may result [16, 19, 20].
@@
To date, the anatomical suitability of TPVR for a given patient is performed by the vendor using a screening CT scan of the RVOT [21]. For example, the screening process of the Harmony valve (Medtronic) was developed as part of the early feasibility study and subsequently refined during subsequent trials based on an engineering analysis that estimated degree of device oversizing (or interference) needed to secure the device in a given RVOT [16, 20, 22]. This analysis is now performed by measuring the circumference of the RVOT at different levels and overlaying the device shape on the perimeter-derived radius of the vessel, resulting in a “perimeter plot” [21]. However, this simplification of the anatomic geometry does not capture the true physical interaction of the device with the complex shape of the RVOT, and fails to account for vessel expansion in the setting of device placement [10]. In addition, many patients are not determined to be candidates for TPVR by standard criteria, but the sensitivity and specificity of these relatively simple screening processes are unknown [21]. As such, patients may be excluded unnecessarily due to conservative guidelines. Adding to the challenge, there are 3 different devices from two different manufacturers currently approved in the United States and over 15 different TPVs across a range of designs and sizes in use or in development internationally [16, 17, 18]. As the number of available devices for TPVR from a range of vendors increases, there is a growing need for an accurate method for assessing patient candidacy and matching the optimal device to an individual patient. In this setting, a physically realistic patient-specific method to select the optimal device for an individual patient within a highly heterogeneous population is of increasing relevance.
@@
In this study, we demonstrate the creation and initial application of an integrated workflow for the finite element (FE) simulation of self-expanding TPV deployment in image-derived, patient-specific RVOT geometries, with the long-term goal of informing patient candidacy and selection of the optimal device for an individual patient. We implemented this open-source capability in the SlicerHeart extension for 3D Slicer and FEBio [23, 24, 25]. Specific contributions include the demonstration of rapid machine learning-based segmentation of the RVOT and coronary arteries, a novel custom interface for TPV selection and positioning within the RVOT, and direct export from SlicerHeart into FEBio for simulation. We then demonstrate virtual TPV implantation of multiple devices in a range of RVOT morphologies and demonstrate visualization and quantification of device shape, compression, and vessel contact to inform optimal device placement in a given patient. Further development of this workflow could inform selection of the optimal TPVR in an individual patient and can be generalizable to any self-expanding medical device.
3 MATERIALS AND METHODS
@@
The modules and other code used for image segmentation are included in the SlicerHeart extension for 3D Slicer and are also available at www.github.com/SlicerHeart [23, 24]. The open-source finite element solver, FEBio version 4.7.0, was used for all FE simulations of TPV deployment [25]. The modeling workflow utilized is displayed in Figure 1.
@@ 3.2 Image Acquisition
A retrospective ECG-gated dual-source scanner (Siemens Healthineers, Forchheim, Germany) was used for image acquisition, with 2 x 128 x 0.6 mm slice collimation. 350 mg/mL Omnipaque (iohexol, GE Healthcare, Chicago, IL), a low-osmolar iodinated contrast, was injected via peripheral intravenous access at a dose of 2 mL/kg, up to 100 mL. The contrast was followed by an injection of normal saline using a dual-head power injector (Medrad Inc, Warrendale, PA). CTA images were imported into 3D Slicer (www.slicer.org) in Digital Imaging and Communications in Medicine (DICOM) format. CT images of three unconstrained TPV devices were acquired on the CT scanner to assist in the creation of realistic device models. In addition, measurements and dimensions were confirmed with Neiko 01407A digital calipers sensitive to 0.01 mm (Zhejiang Kangle Group, Wenzhou, China).
@@ 3.3 RVOT Segmentation
Initial segmentations were generated using threshold paint with 3D sphere brushes and manual revision using the Segment Editor module in 3D Slicer as previously described [10]. The sphere brushes were given a diameter of 2% of the screen size for changes that span multiple slices and 1% for finer edits on a single slice view. A deep learning-based model was trained to automatically segment the RVOT and coronary arteries, as previously described. MONAI’s DeepEdit was modified for this population [26] as we recently described [27]. During model training, the CT images went through a number of transforms for normalization and data augmentation to introduce more variation. All images were resampled to a resolution of 1mm x 1mm x 1mm. During preliminary observations, the intensity values of RVOT and coronary arteries were located in range 125-550 HU. The image intensities were thresholded to 125 $\leq$ I $\leq$ 550 HU and scaled to 0.0 $\leq$ I $\leq$ 1.0, with I representing the image intensity. Data augmentation included random flips, intensity shifts, and 90-degree rotations. Finally, datasets were centrally cropped/padded to have a spatial size of 192 x 192 x 192.
Training was performed on 80 datasets for 50 epochs. As training advanced, the deep-learning model was utilized for the creation of further training samples with manual inspection and, if necessary, modification by a physician with prior segmentation experience (YBC). 100 models were created in total. Segmentations of the RVOT examples herein were made with deep learning and the MONAILabel module in 3D Slicer [28]. The segmentations were subsequently cleaned and smoothed in the Segment Editor module. RVOT segmentations were converted to models and the ends were uncapped using the Dynamic Modeler tool to form hollow models. The automatic segmentation process is shown in Video-1. Figure 2 shows the segmented RVOTs used for this study.
@@ 3.4.1 Device Design
A custom TPVR Simulation module was developed in SlicerHeart to allow placement of a few widely available, commercial TPV devices into the RVOT [24]. The module currently supports positioning of the Harmony TPV-22 (Medtronic, Minneapolis, MN), Harmony TPV-25 (Medtronic, Minneapolis, MN), and the Alterra adaptive prestent (Edwards Lifesciences, Irvine, CA). Models of these devices were created in the Computer Aided Design software, Fusion 360 (Autodesk, San Francisco, CA), from CT imaging and measurements specified in the device documentation [29, 30]. All three TPV devices are made of Nitinol with shape-memory capabilities. The TPV-22 and TPV-25 wireframe devices were modeled with a wire diameter of 0.375 mm. The true Alterra device consists of a laser-cut frame; each wire-like piece has a rectangular cross-section with rounded edges and approximate dimensions of 0.25 mm by 0.32 mm. We modeled the Alterra with a round wire of diameter 0.32 mm for simplicity. We based this value on the diameter of a circle that yields the same cross-sectional area as the actual rectangular cross-section of the Alterra. The TPVs were remeshed as tetrahedral elements using TetWild [31] and GMSH [32], then integrated directly into the TPVR Simulation module as .vtk files for easy user selection. Figure 2 also shows the TPV devices used in this study.
@@ 3.4.2 Device Positioning in Right Ventricular Outflow Tract
Figure 3 shows the procedure for device placement and the TPVR Simulation Module functionality. The TPVR Simulation module uses the vessel centerline and the central axis of the device to position the TPV in the RVOT. The vessel centerline was defined automatically using the Vascular Modeling Toolkit extension centerline extraction functionality or placed manually using an open curve in the Markups module [33] in 3D Slicer. Once a device has been selected, the device is snapped to the centerline for precise placement. If the device imports upside-down with respect to the RVOT model and associated CT image, there is a button to flip the device into the correct orientation while keeping it aligned with the centerline. The user can slide the TPV along the centerline using a slider control or by inputting a value representing the desired location along the centerline. The centerline has been discretized into sections with values from 0 to 100, with 0 corresponding to the most distal point and 100 representing the proximal endpoint. As the device is slid along the vessel centerline, the angle of the device automatically adjusts so that its vertical axis is tangent to the vessel centerline at the center of mass of the TPV. In addition, the TPV can be rotated or translated freely to refine pre-simulation placement. This can be done via menu sliders controlling orientation and spatial position or by using the 3D interactive handler and dragging the device. Should the user wish to change their device type or centerline at any point in the process, they can do so by returning to the Device Selection or Device Positioning subsection, respectively, and selecting a different one. Doing so will not alter the location of the TPV unless they then choose to move it. The device positioning and selection process is highlighted in Video-1.
@@ 3.4.3 Tube Creation for Simulation
To simulate uniform compression of the device prior to TPV expansion, we created a curved tube surface that surrounds the device using the vessel centerline and the MarkupsToModel extension, leveraging vtkTubeFilter [34]. The centerline curve points for the tube are interpolated using global least squares polynomial interpolation. The Centerline Parameters section of the module enables the user to toggle the visibility of the tube and set the radius and number of sides of the tube. The module accommodates radii values of 0 to 50 mm. We chose a radius for the tube that was sufficient to contain the device at the desired position and angle without penetrating the tube, usually 28 mm. The tube’s mesh density depends on the number of points used to form the centerline and the number of sides. Increasing the number of sides adds more elements circumferentially. The number of elements spanning the tube vertically can be increased by resampling the centerline to have more points under Markups. Tube mesh density can be altered to prevent penetration of the device during simulations. Video-1 also highlights the tube creation process.
@@ 3.4.5 FEM Mesh Creation
For each patient case, the shell mesh representing the inner surface of the RVOT was exported from 3D Slicer as a stereolithography (.STL) file to be remeshed as solid hexahedron elements in Hypermesh (Altair, Troy, MI). These surfaces were reduced in density, meshed as quads with an average element size of 1.5 mm, then extruded outward to have a thickness of 1.5 mm [35, 36]. The RVOT mesh was divided into four layers to replicate the orthotropic nature of the artery based on the results of a mesh convergence study observing strain in increasing numbers of layers. The final solid vessel geometries had between 24,000 and 32,000 elements, depending on the size of the vessel. The solid vessel models were saved as Abaqus input .inp files and imported into FEBio. The previously meshed TPV device .vtk and the tube .STL were brought into FEBio as well. The tube was converted to an editable mesh and assigned a shell thickness of 2.0 mm.

### v3

@@ 1 Abbreviations
2D Two
3D Three Dimensional
CHD Congenital Heart Disease
CTA Computed Tomography Angiography
DICOM Digital Imaging and Communications in Medicine
FE Finite Element
HGO Holzapfel Gasser Ogden
RVOT Right Ventricular Outflow Tract
RV-PA Right Ventricle to Pulmonary Artery
STL Stereolithography
ToF Tetralogy of Fallot
TPV Transcatheter Pulmonary Valve
TPVR Transcatheter Pulmonary Valve Replacement
@@ 2 Introduction
Tetralogy of Fallot (ToF) is a form of congenital heart disease (CHD) characterized by pulmonary valve stenosis, right ventricular hypertrophy, a ventricular septal defect, and an overriding aorta. ToF is the most common form of cyanotic CHD, with a prevalence of about 5 out of every 10,000 live births [stallings_national_2024]. In most cases of ToF, a surgical patch is placed to enlarge the pulmonary valve annulus (transannular patch). This patch relieves the obstruction to blood flow in infancy but results in loss of competence of the pulmonary valve. While initially tolerated, over time, severe regurgitation of the pulmonary valve results in right ventricular dilation, ventricular failure, and life-threatening arrhythmias [geva_preoperative_2018, bokma_improved_2023, bokma_propensity_2018, valente_contemporary_2014]. Surgical implantation of a bioprosthetic valve has been used to remedy pulmonary insufficiency and related complications, but has the associated morbidity of open heart surgery on cardiopulmonary bypass [law_transcatheter_2021, gillespie_1-year_2023, crago_pediatric_2023, gartenberg_transcatheter_2023, jolley_toward_2019].
@@
Over the last two decades, transcatheter pulmonary valve replacement (TPVR) has evolved to avoid the risks of open-heart surgery [gartenberg_transcatheter_2023, jolley_toward_2019, kenny_current_2017]. The first transcatheter pulmonary valve (TPV) devices available, the Melody valve (Medtronic, Minneapolis, MN) and Sapien valve (Edwards Lifesciences, Irvine, CA), were balloon-expandable bioprosthetic valves approved only for implantation in the small percentage of patients who received surgically placed cylindrical right ventricle-to-pulmonary artery (RV-PA) conduits as part of their initial repair [gartenberg_transcatheter_2023, jolley_toward_2019, patel_transcatheter_2022]. However, the vast majority of patients with repaired ToF have surgically patched native outflow tracts, which demonstrate significant anatomic variability [schievano_variations_2007, odemis_3d_2025]. As such, nearly 80% of patients with ToF are not candidates for traditional balloon expandable TPV therapies [gillespie_1-year_2023, gartenberg_transcatheter_2023, jolley_toward_2019, kenny_current_2017, patel_transcatheter_2022, schievano_variations_2007, capelli_patient-specific_2010].
@@
In order to meet this need, self-expanding TPV systems that can conform to a wide range of heterogeneous native RVOTs have been developed. In contrast to balloon-expandable TPVs, self-expanding TPVs consist of a spring-like Nitinol frame enclosing a tissue valve [benson_three-year_2020, zahn_first_2018, jin_five-year_2024], which can expand to conform to the vessel and facilitate device retention. However, precise fit of the device in a given individual is critical; if the device is dislodged, it can catastrophically occlude the pulmonary outflow. Further, if the proximal and distal regions of the device do not effectively seal the pulmonary artery, a paravalvular leak or device thrombosis may result [benson_three-year_2020, schoonbeek_implantation_2016, gillespie_patient_2017].
@@
To date, the anatomical suitability of TPVR for a given patient is performed by the vendor using a screening CT scan of the RVOT [mcelhinney_transcatheter_2024]. For example, the screening process of the Harmony valve (Medtronic) was developed as part of the early feasibility study and subsequently refined during subsequent trials based on an engineering analysis that estimated degree of device oversizing (or interference) needed to secure the device in a given RVOT [benson_three-year_2020, gillespie_patient_2017, bergersen_harmony_2017]. This analysis is now performed by measuring the circumference of the RVOT at different levels and overlaying the device shape on the perimeter-derived radius of the vessel, resulting in a “perimeter plot” [mcelhinney_transcatheter_2024]. However, this simplification of the anatomic geometry does not capture the true physical interaction of the device with the complex shape of the RVOT, and fails to account for vessel expansion in the setting of device placement [jolley_toward_2019]. In addition, many patients are not determined to be candidates for TPVR by standard criteria, but the sensitivity and specificity of these relatively simple screening processes are unknown [mcelhinney_transcatheter_2024]. As such, patients may be excluded unnecessarily due to conservative guidelines. Adding to the challenge, there are 3 different devices from two different manufacturers currently approved in the United States and over 15 different TPVs across a range of designs and sizes in use or in development internationally [benson_three-year_2020, zahn_first_2018, jin_five-year_2024]. As the number of available devices for TPVR from a range of vendors increases, there is a growing need for an accurate method for assessing patient candidacy and matching the optimal device to an individual patient. In this setting, a more physically representative, patient-specific geometrical compatibility method to select the optimal device for an individual patient within a highly heterogeneous population is of increasing relevance.
@@
In this study, we demonstrate the creation and initial application of an integrated workflow for the finite element (FE) simulation of self-expanding TPV deployment in image-derived, patient-specific RVOT geometries, with the long-term goal of informing patient candidacy and selection of the optimal device for an individual patient. We implemented this open-source capability in the SlicerHeart extension for 3D Slicer and FEBio [fedorov_3d_2012, lasso_slicerheart_2022, maas_febio_2012]. Specific contributions include the demonstration of rapid machine learning-based segmentation of the RVOT and coronary arteries, a novel custom interface for TPV selection and positioning within the RVOT, and direct export from SlicerHeart into FEBio for simulation. We then demonstrate virtual TPV implantation of multiple devices in a range of RVOT morphologies and demonstrate visualization and quantification of device shape, compression, and vessel contact to inform optimal device placement in a given patient. Further development of this workflow could inform selection of the optimal TPVR in an individual patient and can be generalizable to any self-expanding medical device.
3 Methods
@@
The modules and other code used for image segmentation are included in the SlicerHeart extension for 3D Slicer and are also available at www.github.com/SlicerHeart [fedorov_3d_2012, lasso_slicerheart_2022]. The open-source finite element solver, FEBio version 4.7.0, was used for all FE simulations of TPV deployment [maas_febio_2012]. The modeling workflow utilized is displayed in Figure 1.
@@ 3.2 Image Acquisition
A retrospective ECG-gated dual-source scanner (Siemens Healthineers, Forchheim, Germany) was used for image acquisition, with 2 x 128 x 0.6 mm slice collimation. 350 mg/mL Omnipaque (iohexol, GE Healthcare, Chicago, IL), a low-osmolar iodinated contrast, was injected via peripheral intravenous access at a dose of 2 mL/kg, up to 100 mL. The contrast was followed by an injection of normal saline using a dual-head power injector (Medrad Inc, Warrendale, PA). CTA images were available only for the diastolic phase for all three patients explored in this study. CTA images were imported into 3D Slicer (www.slicer.org) in Digital Imaging and Communications in Medicine (DICOM) format. CT images of three unconstrained TPV devices were acquired on the CT scanner to assist in the creation of realistic device models. In addition, measurements and dimensions were confirmed with Neiko 01407A digital calipers sensitive to 0.01 mm (Zhejiang Kangle Group, Wenzhou, China).
@@ 3.3 RVOT Segmentation
Initial segmentations were generated using threshold paint with 3D sphere brushes and manual revision using the Segment Editor module in 3D Slicer as previously described [jolley_toward_2019]. The sphere brushes were given a diameter of 2% of the screen size for changes that span multiple slices and 1% for finer edits on a single slice view. A deep learning-based model was trained to automatically segment the RVOT and coronary arteries, as previously described. MONAI’s DeepEdit was modified for this population [nguyen_deepedit_2022] as we recently described [barak-corren_image-derived_2025]. During model training, the CT images went through a number of transforms for normalization and data augmentation to introduce more variation. All images were resampled to a resolution of 1mm x 1mm x 1mm. During preliminary observations, the intensity values of RVOT and coronary arteries were located in range 125-550 HU. The image intensities were thresholded to 125 $\leq$ I $\leq$ 550 HU and scaled to 0.0 $\leq$ I $\leq$ 1.0, with I representing the image intensity. Data augmentation included random flips, intensity shifts, and 90-degree rotations. Finally, datasets were centrally cropped/padded to have a spatial size of 192 x 192 x 192.
Training was performed on 80 datasets for 50 epochs. As training advanced, the deep-learning model was utilized for the creation of further training samples with manual inspection and, if necessary, modification by a physician with prior segmentation experience (YBC). 100 models were created in total. Segmentations of the RVOT examples herein were made with deep learning and the MONAILabel module in 3D Slicer [diaz-pinto_monai_2024]. The segmentations were subsequently cleaned and smoothed in the Segment Editor module. RVOT segmentations were converted to models and the ends were uncapped using the Dynamic Modeler tool to form hollow models. The automatic segmentation process is shown in Video-1. Figure 2 shows the segmented RVOTs used for this study. The respective RVOT morphology types based on perimeter-derived diameters at different lengths along the RVOT [schievano_variations_2007] are also highlighted in 2.
@@ 3.4.1 Device Design
A custom TPVR Simulation module was developed in SlicerHeart to allow placement of a few widely available, commercial TPV devices into the RVOT [lasso_slicerheart_2022]. The module currently supports positioning of the Harmony TPV-22 (Medtronic, Minneapolis, MN), Harmony TPV-25 (Medtronic, Minneapolis, MN), and the Alterra adaptive prestent (Edwards Lifesciences, Irvine, CA). Models of these devices were created in the Computer Aided Design software, Fusion 360 (Autodesk, San Francisco, CA), from CT imaging and measurements specified in the device documentation [medtronic_tpv_2021, edwards_lifesciences_edwards_2021]. All three TPV devices are made of Nitinol with shape-memory capabilities. The TPV-22 and TPV-25 wireframe devices were modeled with a wire diameter of 0.375 mm. The true Alterra device consists of a laser-cut frame; each wire-like piece has a rectangular cross-section with rounded edges and approximate dimensions of 0.25 mm by 0.32 mm. We modeled the Alterra with a round wire of diameter 0.32 mm for simplicity. We based this value on the diameter of a circle that yields the same cross-sectional area as the actual rectangular cross-section of the Alterra. The TPVs were remeshed as tetrahedral elements using TetWild [hu_tetrahedral_2018] and GMSH [geuzaine_gmsh_2009], then integrated directly into the TPVR Simulation module as .vtk files for easy user selection. Figure 2 also shows the TPV devices used in this study.
@@ 3.4.2 Device Positioning in Right Ventricular Outflow Tract
Figure 3 shows the procedure for device placement and the TPVR Simulation Module functionality. The TPVR Simulation module uses the vessel centerline and the central axis of the device to position the TPV in the RVOT. The vessel centerline was defined automatically using the Vascular Modeling Toolkit extension centerline extraction functionality or placed manually using an open curve in the Markups module [piccinelli_framework_2009] in 3D Slicer. Once a device has been selected, the device is snapped to the centerline for precise placement. If the device imports upside-down with respect to the RVOT model and associated CT image, there is a button to flip the device into the correct orientation while keeping it aligned with the centerline. The user can slide the TPV along the centerline using a slider control or by inputting a value representing the desired location along the centerline. The centerline has been discretized into sections with values from 0 to 100, with 0 corresponding to the most distal point and 100 representing the proximal endpoint. As the device is slid along the vessel centerline, the angle of the device automatically adjusts so that its vertical axis is tangent to the vessel centerline at the center of mass of the TPV. In addition, the TPV can be rotated or translated freely to refine pre-simulation placement. This can be done via menu sliders controlling orientation and spatial position or by using the 3D interactive handler and dragging the device. Should the user wish to change their device type or centerline at any point in the process, they can do so by returning to the Device Selection or Device Positioning subsection, respectively, and selecting a different one. Doing so will not alter the location of the TPV unless they then choose to move it. The device positioning and selection process is highlighted in Video-1.
@@ 3.4.3 Tube Creation for Simulation
To simulate uniform compression of the device prior to TPV expansion, we created a curved tube surface that surrounds the device using the vessel centerline and the MarkupsToModel extension, leveraging vtkTubeFilter [schroeder_visualization_2006]. The centerline curve points for the tube are interpolated using global least squares polynomial interpolation. The Centerline Parameters section of the module enables the user to toggle the visibility of the tube and set the radius and number of sides of the tube. The module accommodates radii values of 0 to 50 mm. We chose a radius for the tube that was sufficient to contain the device at the desired position and angle without penetrating the tube, usually 28 mm. The tube’s mesh density depends on the number of points used to form the centerline and the number of sides. Increasing the number of sides adds more elements circumferentially. The number of elements spanning the tube vertically can be increased by resampling the centerline to have more points under Markups. Tube mesh density can be altered to prevent penetration of the device during simulations. Video-1 also highlights the tube creation process.
@@ 3.4.5 FEM Mesh Creation
For each patient case, the shell mesh representing the inner surface of the RVOT was exported from 3D Slicer as a stereolithography (.STL) file to be remeshed as quad elements in Blender (V4.4). These surfaces were reduced in density and meshed as quads with an average element size of 1.5 mm. This quad meshed RVOT surface was saved as a .ply format from Blender and subsequently imported into FEBio. The quad mesh was then extruded normally outward to have a thickness of 1.5 mm [donahue_finite_2024, vanderveken_mechano-biological_2020]. The RVOT mesh was divided into four layers to replicate the orthotropic nature of the artery based on the results of a mesh convergence study observing strain in increasing numbers of layers. The final solid vessel geometries had between 24,000 and 32,000 elements, depending on the size of the vessel. The previously meshed TPV device .vtk and the tube .STL were brought into FEBio as well. The tube was converted to an editable mesh and assigned a shell thickness of 2.0 mm.